OpenStudy (anonymous):

What is the value of w? w – 13 = 72 w =

OpenStudy (anonymous):

@luckycoins888

OpenStudy (anonymous):

What is the value of w? w – 13 = 72 w =

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (shadowlegendx):

w - 13 = 72 Add 13 to both sides. w = 85

OpenStudy (anonymous):

help ppplllzzzz!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the value of h? 57 + h = 81 h =

OpenStudy (anonymous):

...

OpenStudy (anonymous):

25?

OpenStudy (luckycoins888):

24